OpenStudy (anonymous):

Can you check my solution? Question: For 0≤t≤5, find when t(cos t)>0. My answer: t>cos^-1(0). Thanks.

OpenStudy (anonymous):

0<cost<=1 all 0<=t<=5

OpenStudy (anonymous):

thus tcost>0 all 0<t<=5 except t=0

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

for \(\frac{\pi}{2}<t<\frac{3\pi}{2}\) we have \(\cos(t)<0\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so \(t\cos(t)<0\) on that interval as well

OpenStudy (anonymous):

point is that when \(0<t<5\) the solution to \(t\cos(t)>0\) is the same as the solution to \(\cos(t)>0\)

OpenStudy (anonymous):

@chanh_chung question asks \(>0\) not \(>1\) i think

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73 would the right answer be 0<t<pi/2 and t>3(pi)/2

OpenStudy (anonymous):

I say tcost>0 all 0<t<=5 except t=0 because t=0 is tcost=0